OpenStudy (anonymous):

\[4b^{2} + 8b = 0\] Please show me the steps to solve

OpenStudy (anonymous):

Factor out b would be your first step

OpenStudy (anonymous):

b(b+8)=0 so now you can say b=0 and b+8=0

OpenStudy (anonymous):

usually when factoring out, you subtract from both sides? here you just eliminate 4?

OpenStudy (anonymous):

Wow ok got my glasses now lol. you would factor out 4b 4b(b+2)=0

OpenStudy (anonymous):

XD I was about to say

OpenStudy (anonymous):

Yeah sorry about that. so 4b=0 b=0 and b+2=0 b=-2

OpenStudy (anonymous):

What happens to the 8? did you subtract 4 from it as well?

OpenStudy (anonymous):

When you pull out the 4b its like dividing every term by 4b 8b/4b=2

OpenStudy (anonymous):

I see so you would distribute divide 4b to 4b,8b and 0.

OpenStudy (anonymous):

Can you rewrite it out again?

OpenStudy (anonymous):

Im not sure what you're not understanding Factor out 4b, which divides every term by 4b 4b(4b^2/4b+8b/4b) 4b(b+2)